5.2.5 Question Help Assume the random variable x is normally distributed with mean H 86 and standard deviation ơ-4. Find the indicated probability P(75<x 80) Round to four decimal places as needed.)

Assume the random variable \(\mathrm{x}\) is normally distributed with mean \(\mu=86\) and standard deviation \(\sigma=4\). Find the indicated probability.

$$ \begin{array}{r} P(75<x<80) \\ P(75<x<80)= \end{array} $$

(Round to four decimal places as needed.)
Answer #1

Solution:- P( 75 < x < 80) = 0.0638

Mean = 86, S.D = 4

x1 = 75

x2 = 80

By applying normal distribution:-

z = \frac{x-\mu }{\sigma }

z1 = - 2.75

z2 = -1.50

P( -2.75 < z < -1.50) = P(z > -2.75) - P(z > -1.50)

P( -2.75 < z < -1.50) = 0.997 - 0.9332

P( -2.75 < z < -1.50) = 0.0638